OpenStudy (anonymous):

What will the exponent be when you simplify the following problem? (Make sure you type just the value of the exponent only. Do not skip any spaces.) x^5/x^2=x^?

OpenStudy (anonymous):

whenever you divide exponents, you subtract them johana:)

OpenStudy (nurali):

\[x^{5-2}\]

OpenStudy (anonymous):

\[\frac{ x^5 }{ x^2 }=x^5-^{2}\] thats the best i can get it it wont lent me put a subtraction sign between the exponents